# ACE: Abstractions for Communicating Efficiently

### v1

@@ 3.4 Full Algorithm
Algorithm 1 shows how the components of ACE integrate. The algorithm alternates between the Communication Phase and the Abstraction Phase for $s$ steps, where $s$ is determined experimentally [^1] We found that 20 steps were a good choice for our setting, after which the language stabilises and no more abstractions are introduced in the emergent language..
@@
The Communication Phase is repeated for $e$ iterations, determined experimentally. In each iteration, programs are sampled from $\pi_{arch}$ by the bandit’s $\epsilon$-greedy strategy and are used to form a dataset of grid-state transitions. This dataset forms the basis of our signalling game and is used to train $\pi_{comm}$, $\pi_{bldr}$ and to update the $Q$-values of actions. Here, we expect a language to emerge which is subject to pressures for efficient communication.
@@
In the Abstraction Phase sub-sequences in the programs are identified. The sub-sequence that maximises Equation 1 is selected. This abstraction, $a_{new}$, is introduced into $\cal A$. $a_{new}$ may allow existing programs to be rewritten: in this case the resulting shorter programs are introduced into $\pi_{arch}$ for the respective goal-scene. Finally, $Q(a_{new})$ is initialised. More detailed descriptions of ACE’s phases can be found in Appendix A.2.
Algorithm 1 Abstractions for Communicating Efficiently (ACE)
@@
8:        Play signalling games with $D$ to train $\pi_{comm}$, $\pi_{bldr}$ and $Q$-values using Eq. 3, 4 and 5
@@
4 Empirical Evaluation
To evaluate ACE we investigate the following questions: 1) Does ACE result in a final language allowing for shorter communications? 2) Does communicative pressures impact what abstractions get adopted?
The behaviour of ACE is compared to two naïve baselines. The first, No abstractions, is ACE without the abstraction phase. The second, Greedy, always picks the shortest program. We use reward (also referred to as reconstruction accuracy) and average program length to compare these approaches. Unless otherwise stated results obtained are averaged over $16$ runs. We provide further implementation details in Appendix A.3.
Figure 2: Average program length vs. steps for ACE, Greedy and No abstractions. Line indicates mean performance and shaded regions indicate the 95% confidence interval.
Figure 3: Test reward vs. steps for ACE, Greedy and No abstractions. Line indicates mean performance and shaded regions indicate the 95% confidence interval.
4.1 ACE reduces program length
ACE does indeed converge to shorter programs over interaction steps, as seen in Figure 2. When we compare ACE to the No abstraction variant, the program lengths are halved, going from an average length of $10$ to one of $4.92\pm 0.20$, indicating that a more efficient language has been derived. The Greedy variant reduces program lengths even more, however, this comes at a price. In Figure 3 we see drastic drops in accuracy for Greedy every time an abstraction is introduced, whereas ACE’s slower introduction leads to smaller drops in accuracy. For a while, both strategies manage to recover near full accuracy, but after some time, the Greedy communication fails to recover and becomes progressively worse.
4.2 The Communicative Pressures on Abstractions
Figure 4: The abstractions introduced by ACE in an example run. We circle in green the abstractions that are adopted by the architect in its final language.
(a) The first abstraction, resembling a tower, being introduced successfully.
(b) The thirteenth abstraction, resembling an X, is introduced, but not retained.
Figure 5: Different abstractions being introduced. We show the mean and 95% confidence interval for Q-value (black) and frequency (blue) versus epochs (of which there are $40$ in a step) over $20$ random seeds.
Figure 4 shows the abstractions that ACE introduces in an example run [^2] Further examples, which are quite similar, can be found in our code repository.. Abstractions circled in green are in the final language. ACE is able to compose abstractions: smaller abstractions may be utilised later in larger abstractions. For example, abstraction 17 is a sequence of two instances of abstraction 3.
We note that not all of the 20 proposed abstractions are retained in the final language. In Figure 5(a) and 5(b), we show abstractions that experience different fates. Figure 5(a), shows the Q-value and frequency of abstraction 1 from Figure 4 (a tower), which was retained. The tower is very common in our dataset, appearing $1380$ times. Figure 5(b), shows the same for abstraction 13 from Figure 4 (the letter X), which was not retained. The X is relatively uncommon and appears only 60 times. In the case of the tower, after a period of exploration, ACE readily adopts it. We note a brief period of instability that occurs at step $2$, while a second abstraction is being introduced, after which the usage of the tower gradually stabilises. The X on the other hand, has a small spike in frequency initially, but it quickly drops to zero as the agents abandon it, having found it too hard to learn.
Figure 7: Two goal scenes with representations of their initial and final programs chosen by ACE.
5 Discussion
Addressing the challenge in (Ho et al. 2019), we initiated the study of how abstractions could be used to make communication between artificial agents more efficient. In contrast to (McCarthy et al. 2021) who studied this question with human subjects, ours is the first to explore it in the setting of artificial agents.
We have demonstrated that ACE exhibits the capability of developing a compact collaborative language, resulting in shorter programs. The size of the language is the consequence of pressures that naturally arise through the need to communicate about a shared task. This is achieved by ACE through a novel combination of EC, library learning and bandits. Our work serves as a bridge between abstraction learning and efficient communication (Kemp and Regier 2012; Gibson et al. 2017; Zaslavsky et al. 2019; Gibson et al. 2019; Denić and Szymanik 2024). It builds on previous research showing that Emergent Communication can be used to learn efficient communication in specific domains, such as numerals (Carlsson, Dubhashi, and Johansson 2021) and colors (Carlsson, Dubhashi, and Regier 2023), and extends this to the learning of abstractions.
Beyond importance within cognitive science, we believe our results have real implications for machine learning. Significant research effort is being invested into exploring the role of natural language as a mechanism for humans to provide instructions to intelligent agents (Brohan et al. 2023; Shi et al. 2024). It is natural for humans to introduce abstractions and we should create systems that can accommodate this. Research promoting this capability can facilitate improved cooperation between humans and agents. We believe that ACE represents a step towards equipping intelligent agents with flexible and extendable languages.
Limitations
So far, we have not applied ACE to domains beyond the Architect-Builder game, and extension towards other settings, such as the ARC challenge (Chollet 2019) or PizzaCommonSense (Diallo et al. 2024) is left to future work. In novel environments such as these, our assumption that a set of initial programs exists for each goal-state may not be realisable. This would require also initial programs to be learned which would further complicate the task of learning to communicate. Furthermore, our work does not yet analyse ACE’s languages from an information-theoretic perspective, e.g. as a trade-off between average program length and vocabulary size, situating them in relation to a Pareto frontier.
Future Work
ACE presents a step towards understanding how abstractions are introduced for instruction sequences. It has been shown by Krauss and Weinheimer (1964) that abstractions also appear within repeated exchanges expressed in natural language. This is a complementary perspective on the role of abstractions within human communication. We intend to integrate both of these perspectives into a unified model using LLM-based agents like those in (Brohan et al. 2023).
6 Related Literature
Emergent communication in the Architect-Builder game was first studied in humans by McCarthy et al. (2021), however only on a smaller dataset with only 12 goal-scenes. They also propose a simple Bayesian model for the task, inspired by Goodman and Frank (2016) and Ellis et al. (2021). However, this is not a proper multi-agent system and is limited to assuming that abstractions are introduced automatically by a centralised process in both the architect and builder at the same time. Our work addresses this major limitation by situating the task in a multi-agent system where the introduction of abstractions is controlled by the architect, a more plausible representation. When a new abstraction is introduced, ACE allows the architect and builder to negotiate a change to their common language to incorporate it, or if it proves difficult to learn, discard it. Furthermore, our setting does not require an explicit parameter to keep the language size manageable, this appears automatically as a consequence of communicative pressures in ACE. Jergéus et al. (2022) conduct a preliminary study into using multi-agent reinforcement learning for the Architect-Builder task, but make many simplifications. Most notably, there is no learning of a messaging policy and the abstractions are assumed to be immediately understood by the builder. Under this assumption, this actually collapses to a single agent setting and is not subject to pressures for efficient communication.
While EC has been used to study areas such as the impact of populations (Chaabouni et al. 2022; Kim and Oh 2021; Rita et al. 2022), compositionality (Mordatch and Abbeel 2018; Chaabouni et al. 2020; Ren et al. 2020), and naming conventions in semantic categories (Carlsson, Dubhashi, and Regier 2023; Carlsson, Dubhashi, and Johansson 2021), we introduce it as a tool to study the process of abstraction and combine it with library learning and bandits.
7 Conclusions
When human conversational partners repeatedly cooperate on a task, they tend to develop more concise language. In this paper, we study this process within a collaborative building game known as the architect-builder game. We propose a mechanistic explanation that integrates bandit models with combinatorial actions, emergent communication, and library learning to model the abstraction process. Together, these components enable for the proposal and evaluation of new abstractions based on their communicative efficiency and their impact on program length. We find that our method aligns with human tendencies towards more concise language over repeated interactions. In comparison to previous approaches, our model represents a more naturalistic approach, which removes restrictions on the manner in which abstractions are introduced and naturally converges to a fixed policy as a result of emergent communication thereby reducing dependence on hard-coded parameters.
Appendix A Appendices
A.1 Dataset
Figure 8: The $31$ shapes that in pair compose each goal-scene.
Each goal-scene in the dataset is composed of two shapes placed next to each other, as in Figure 1. These are sampled from 31 shapes we design and report in Figure 8. Each shape is meant to resemble an uppercase or lowercase letter from the English alphabet. We note that there are multiple subshapes that recur several times within the 31 shapes, making our dataset a prime candidate for studying the introduction of abstraction.
A.2 Algorithms composing ACE
@@
In this Section, we provide more detailed algorithms for the focal points of ACE’s methodology. Algorithms 2 and 3 refer to the Communication Phase, which we break down further in two parts: the dataset generation for the signalling game and the communication round done via EC. Algorithm 4 instead refers to the Abstraction Phase.
Algorithm 2 Dataset generation via bandit
@@
4:     Compute Q-values for programs (Equation 6)
@@
Algorithm 3 A round of communication of ACE
@@
Algorithm 4 Abstraction phase
@@
4:     Compute Q-values for programs (Equation 6)
@@
A.3 Implementation Details
We implement ACE within Python version 3.9.18 and PyTorch version 2.2.1 (Paszke et al. 2019) and run experiments on Apple M2 Pro CPU (16 GB) on MacOS 14.4.1. The communication policies, $\pi_{comm}$ and $\pi_{bldr}$, are deep neural networks which have $1$ and $2$ hidden layers, respectively. Each hidden layer contains $200$ neurons and utilises ReLU activation functions. The outputed layer of $\pi_{comm}$ is comprised of $30$ neurons, which is its maximum vocabulary size. The output layer of $\pi_{bldr}$ is comprised of $81$ neurons, which is the size of the grid. Each output neuron in $\pi_{bldr}$ represents the builder’s belief that the grid index is occupied. We use the gumbel-softmax relaxation with the hard parameter so that the networks can be differentiated end-to-end. The model is optimised to minimise binary cross entropy error and we use a learning rate of $0.0009$ in conjunction with the ADAM optimiser (Kingma and Ba 2014). The bandit hyperparameters $\alpha$, $\gamma$, $\epsilon$ and $q_{init}$ are set to $0.5$, $0.99$, $0.1$, and $0.0$ respectively. Note, that before an abstraction is introduced we prune all but the $3$ best programs (determined empirically) for each goal-scene in $\pi_{arch}$ to not incur an exponential growth in the number of programs for each goal-scene. We also set the number of epochs $e$ to 40. All hyperparameters chosen are experimentally determined through grid search.
A.4 Q-value initialisation: Pessimism is best
@@
In §3.2, we considered the impact of initialisation on learning within ACE. Here, we establish its impact on learning of the common “tower” abstraction. We compare a pessimistic strategy ($q_{init}=0$) to an optimistic strategy ($q_{init}=1$) in terms of cumulative regret, defined as $Regret_{T}=\sum_{t=1}^{T}\left(r^{*}-r_{t}\right)$. In Figure 9, we see that a pessimistic initialisation of 00 incurs significantly less cumulative regret than an optimistic value of $1$. This is attributable to the optimistic variant’s tendency to force the abstraction into use before the builder has had the opportunity to become familiar with it. The pessimistic agent introduces the new abstraction more gradually while still providing the opportunity for learning of a common language.
@@
Figure 9: Cumulative regret for an optimistic and pessimistic initialisation of Q-values described in Equation 5. Line indicates mean performance and shaded regions indicate the 95% confidence interval for $16$ random seeds. Epoch interval is between step $1$ and step $2$.
Acknowledgments
This work was supported by funding from from the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and the Swedish Research Council. The computations in this work were enabled by resources provided by the Swedish National Infrastructure for Computing (SNIC). We also extend our thanks to Emil Carlsson for insightful discussions at various points throughout the project.

### v2

@@ Acknowledgments
Algorithm 1 shows how the components of PACE integrate. The algorithm alternates between the Communication Phase and the Abstraction Phase for $s$ steps, where $s$ is determined experimentally [^5] We found that 20 steps were a good choice for our setting, after which the language stabilises and no more abstractions are introduced in the emergent language..
@@
The Communication Phase is repeated for $e$ epochs, determined experimentally. At each iteration, programs are sampled from $\pi_{arch}$ (the table of programs for scenes) by the bandit’s $\epsilon$-greedy strategy and are used to form a dataset of grid-state transitions. This dataset forms the basis of our signalling game and is used to train $\pi_{comm}$, $\pi_{bldr}$ and to update the $Q$-values of actions. Here, we expect a language to emerge which is subject to pressures for efficient communication.
@@
In the Abstraction Phase sub-sequences in the programs are identified. The sub-sequence that maximises Equation 1 is selected. This abstraction, $a_{new}$, is introduced into $\cal A$. $a_{new}$ may allow existing programs to be rewritten: in this case, the resulting shorter programs are introduced into $\pi_{arch}$ for the respective goal-scene. Finally, $Q(a_{new})$ is initialised.
@@
8:        Play signalling games with $D$ to train $\pi_{comm}$, $\pi_{bldr}$ and $Q$-values
@@
Algorithm 1 Procedural Abstractions for Communicating Efficiently (PACE)
@@
We also provide more detailed algorithms for the focal points of PACE’s methodology. Algorithms 2 and 3 refer to the Communication Phase, which we break down further into two parts: the dataset generation for the signalling game and the communication round done via EC. Algorithm 4 instead refers to the Abstraction Phase.
@@
4:     Compute Q-values for programs
@@
Algorithm 2 Dataset generation via bandit
@@
Algorithm 3 A round of communication of PACE
@@
4:     Compute Q-values for programs
@@
Algorithm 4 Abstraction phase
Q-value initialisation: Pessimism is best
@@
It is not obvious how the initialisation of the Q-values for new abstractions may impact regret. Compelling cases can be made for both optimistic (a more aggressive introduction) and pessimistic (a more gradual introduction) strategies. Empirically we find the pessimistic strategy works better in our case. Here, we establish its impact on learning of the common “tower” abstraction. We compare a pessimistic strategy ($q_{init}=0$) to an optimistic strategy ($q_{init}=1$) in terms of cumulative regret, defined as $Regret_{T}=\sum_{t=1}^{T}\left(r^{*}-r_{t}\right)$. In Figure 8, we see that a pessimistic initialisation of 00 incurs significantly less cumulative regret than an optimistic value of $1$. This is attributable to the optimistic variant’s tendency to force the abstraction into use before the builder has had the opportunity to become familiar with it. The pessimistic agent introduces the new abstraction more gradually while still providing the opportunity for learning of a common language.
@@
Figure 8: Cumulative regret for an optimistic and pessimistic initialisation of Q-values. Line indicates mean performance and shaded regions indicate the 95% confidence interval for $16$ random seeds. Epoch interval is between step $1$ and step $2$.
Examples of Abstractions Introduced
In Figure 9 we report the abstractions that PACE introduces in a single run. Abstractions circled in green are in the final language.
Figure 9: The abstractions introduced by PACE in an example run. We circle in green the abstractions that are adopted by the architect in its final language.
Implementation Details
We implement PACE within Python version 3.9.18 and PyTorch version 2.2.1 [Paszke et al., 2019] and run experiments on Apple M2 Pro CPU (16 GB) on MacOS 14.4.1. The communication policies, $\pi_{comm}$ and $\pi_{bldr}$, are deep neural networks which have $1$ and $2$ hidden layers, respectively. Each hidden layer contains $200$ neurons and utilises ReLU activation functions. The outputted layer of $\pi_{comm}$ is comprised of $30$ neurons, which is its maximum vocabulary size. The output layer of $\pi_{bldr}$ is comprised of $81$ neurons, which is the size of the grid. Each output neuron in $\pi_{bldr}$ represents the builder’s belief that the grid index is occupied. We use the gumbel-softmax relaxation with the hard parameter so that the networks can be differentiated end-to-end. The model is optimised to minimise binary cross entropy error and we use a learning rate of $0.0009$ in conjunction with the ADAM optimiser [Kingma and Ba, 2014]. The bandit hyperparameters $\alpha$, $\gamma$, $\epsilon$ and $q_{init}$ are set to $0.5$, $0.99$, $0.1$, and $0.0$ respectively. Note, that before an abstraction is introduced we prune all but the $3$ best programs (determined empirically) for each goal-scene in $\pi_{arch}$ to not incur an exponential growth in the number of programs for each goal-scene. We also set the number of epochs $e$ to 40. All hyperparameters chosen are experimentally determined through grid search.
Figure 10: The $31$ shapes that in pair compose each goal-scene.